# Crosscap states and duality of Ising field theory in two dimensions

### v1

@@ Abstract
×$\times\prod_{j=1}^{2n}\left[|w_{j}|^{-2m\epsilon_{j}}\left(1+\frac{1}{w_{j}\bar{w}_{j}}\right)^{-1/4}\right]\prod_{1\leq j<l\leq 2n}\left|\frac{1}{w_{j}-w_{l}}\cdot\left(1+\frac{1}{w_{j}\bar{w}_{l}}\right)\right|^{-\epsilon_{j}\epsilon_{l}/2}\,.$ (S149)
@@
${}_{\mathrm{NS}}\langle 0|\sigma(w_{1},\bar{w}_{1})\sigma(w_{2},\bar{w}_{2})|1\rangle\rangle_{\mathcal{C}}$ $=\frac{1}{|w_{1}-w_{2}|^{1/4}}\sqrt{\frac{1}{2}\left[(1-\eta)^{-1/4}+(1-\eta)^{1/4}\right]}$
@@
${}_{\mathrm{NS}}\langle 0|\sigma(w_{1},\bar{w}_{1})\sigma(w_{2},\bar{w}_{2})|\mathcal{C}_{\pm}\rangle$ $=\sqrt{\frac{2+\sqrt{2}}{2}}\frac{1}{|w_{1}-w_{2}|^{1/4}}\cdot(1-\eta)^{-1/8}\left[\frac{\sqrt{2}}{2}\sqrt{1+\sqrt{1-\eta}}\pm\frac{2-\sqrt{2}}{2}\sqrt{1-\sqrt{1-\eta}}\right]\,,$ (S153)
where $\eta$ is the crosscap cross ratio [Eq. (S138)]. This corresponds to Eq.(12) in the main text, where the correlator with the crosscap state $|\mathcal{C}_{+}\rangle$ is consistent with that obtained using the sewing constraints [22] and the conformal partial wave decomposition [32, 33].
@@
The conformal perturbation theory for the crosscap overlap is given in Eq. (15) in the main text. Here, we provide a simple derivation and then formulate the perturbation expansion for the general crosscap state $|\mathcal{C}\rangle$ up to second order. The application to the perturbed Ising CFT is specifically discussed.
@@
where we expect that the perturbed ground-state energy scales as $E_{0}=-\frac{\pi}{6L}c(s)$ and the energy gap scales as $\Delta E(s,L)\sim\frac{1}{L}$. Here, $c(s)$, with $c(0)=c$, is the “running” central charge, denoting the deformed central charge under the perturbation. As $\beta\to\infty$, the contribution of the excited states decays exponentially, leaving only the ground state contribution. Therefore, in the interaction picture, we can extract $W(s)$ as
$W(s)=\lim_{\beta\to\infty}\left[\ln\langle\psi_{0}(0)|e^{\beta H_{0}}e^{-\beta H}|\psi_{0}(0)\rangle-\frac{\pi\beta}{6L}\delta c(s)\right]=\lim_{\beta\to\infty}\left[\left\langle\mathcal{T}e^{-\int_{0}^{\beta}\mathrm{d}\tau\,H_{1}(\tau)}\right\rangle_{\mathrm{c}}-\frac{\pi\beta}{6L}\delta c(s)\right]\,,$ (S158)
where $\langle\mathcal{T}e^{-\int_{0}^{\beta}\mathrm{d}\tau\,H_{1}(\tau)}\rangle_{\mathrm{c}}$ denotes the connected contribution of $\langle\psi_{0}(0)|\mathcal{T}e^{-\int_{0}^{\beta}\mathrm{d}\tau\,H_{1}(\tau)}|\psi_{0}(0)\rangle$, via the linked cluster theorem: $\ln\langle\psi_{0}(0)|\mathcal{T}e^{-\int_{0}^{\beta}\mathrm{d}\tau\,H_{1}(\tau)}|\psi_{0}(0)\rangle=\langle\mathcal{T}e^{-\int_{0}^{\beta}\mathrm{d}\tau\,H_{1}(\tau)}\rangle_{\mathrm{c}}$, and
@@
is the change of ”runing” central charge.
We can obtain the formal perturbation series by expanding the time-ordered exponential $\mathcal{T}e^{-\int_{0}^{\beta}\mathrm{d}\tau\,H_{1}(\tau)}$ in powers of the coupling $g$. Specifically, for $W(s)$, we expand $\langle\psi_{0}(0)|\mathcal{T}e^{-\int_{0}^{\beta}\mathrm{d}\tau\,H_{1}(\tau)}|\psi_{0}(0)\rangle$ as
@@
where each order of the perturbation series can be split into a divergent term $\frac{2\beta}{6L}C_{2n}$, scaling as $\frac{\beta}{L}$ when $\beta\to\infty$, and a subleading finite term $W_{2n}$. The series expansion of $W(s)$ is obtained as $W(s)=\sum_{n=1}^{\infty}W_{2n}s^{2n}$ and the “running” central charge is obtained as $c(s)=c+\sum_{n=1}^{\infty}C_{2n}s^{2n}$.
@@
The cosscap state $|\mathcal{C}\rangle$ for a CFT can be expressed as a linear combination of the crosscap Ishibashi states $|a\rangle\rangle_{\mathcal{C}}$, which are labeled by the primary fields $\varphi_{a}$ with the conformal weights $h_{\alpha}=\bar{h}_{\alpha}$:
where the coefficient corresponding to the identity operator is chosen to be positive: $A_{1}>0$. The one-point crosscap correlator is given by:
@@
and the two-point crosscap correlator can be determined up to an unknown prefactor $G_{a}(\eta)$ as a function of the crosscap cross ratio [22, 32, 33]:
$\langle\psi_{0}(0)|\varphi_{a}(w_{1},\bar{w}_{1})\varphi_{a}(w_{2},\bar{w}_{2})|\mathcal{C}\rangle=\frac{\mathcal{A}_{1}\cdot G_{a}(\eta)}{|w_{1}-w_{2}|^{4h_{a}}}\,,$ (S163)
@@
with $Z_{0}=\mathcal{A}_{1}$ by definition. Then, the perturbative expansion of the crosscap overlap up to second order is given by
$\langle\psi_{0}(s)|\mathcal{C}\rangle=\mathcal{A}_{1}+Z_{1}\cdot s+\left(Z_{2}+\frac{1}{2}\mathcal{A}_{1}\cdot W_{2}\right)\cdot s^{2}+\mathcal{O}(s^{3})\,.$ (S166)
@@
$Z_{1}\cdot s=-\int_{0}^{\infty}\mathrm{d}\tau\langle\psi_{0}(0)|H_{1}(\tau)|\mathcal{C}\rangle=g\int_{0}^{\infty}\mathrm{d}\tau\int_{0}^{L}\mathrm{d}x\,\langle\psi_{0}(0)|\varphi_{a}(z,\bar{z})|\mathcal{C}\rangle~{}$ (S167)
@@
$Z_{2}\cdot s^{2}$ $=\frac{1}{2}\int_{0}^{\infty}\mathrm{d}\tau_{1}\int_{0}^{\infty}\mathrm{d}\tau_{2}\left[\langle\psi_{0}(0)|H_{1}(\tau_{1})H_{1}(\tau_{2})|\mathcal{C}\rangle-\mathcal{A}_{1}\langle\psi_{0}(0)|H_{1}(\tau_{1})H_{1}(\tau_{2})|\psi_{0}(0)\rangle\right]$
$=\frac{g^{2}}{2}\int_{0}^{\infty}\mathrm{d}\tau_{1}\int_{0}^{L}\mathrm{d}x_{1}\int_{0}^{\infty}\mathrm{d}\tau_{2}\int_{0}^{L}\mathrm{d}x_{2}\left[\langle\psi_{0}(0)|\varphi_{a}(z_{1},\bar{z}_{1})\varphi_{a}(z_{2},\bar{z}_{2})|\mathcal{C}\rangle-\mathcal{A}_{1}\langle\psi_{0}(0)|\varphi_{a}(z_{1},\bar{z}_{1})\varphi_{a}(z_{2},\bar{z}_{2})|\psi_{0}(0)\rangle\right]~{}$ (S168)
@@
$Z_{2}\cdot s^{2}$ $=\frac{g^{2}}{2}\left(\frac{2\pi}{L}\right)^{4h_{a}}\mathcal{A}_{1}\int_{0}^{\infty}\mathrm{d}\tau_{1}\int_{0}^{L}\mathrm{d}x_{1}\int_{0}^{\infty}\mathrm{d}\tau_{2}\int_{0}^{L}\mathrm{d}x_{2}\frac{|w_{1}w_{2}|^{2h_{a}}}{|w_{1}-w_{2}|^{4h_{a}}}\left(G_{a}(\eta)-1\right)$
$=\left[\frac{\mathcal{A}_{1}}{2(2\pi)^{3-4h_{a}}}\int_{0}^{1}\mathrm{d}\rho_{1}\int_{0}^{1}\mathrm{d}\rho_{2}\int_{0}^{2\pi}\mathrm{d}\theta\frac{(\rho_{1}\rho_{2})^{2h_{a}-1}}{|\rho_{1}-\rho_{2}e^{i\theta}|^{4h_{a}}}\left(G_{a}(\eta)-1\right)\right]\cdot s^{2}\,,$ (S172)
where the new integral variables read $(\rho_{1},\rho_{2},\theta)=(e^{-\frac{2\pi}{L}\tau_{1}},e^{-\frac{2\pi}{L}\tau_{2}},\frac{2\pi(x_{1}-x_{2})}{L})$. In the above derivation, the one- and two-point crosscap correlators [Eqs. (S162) and  (S163)] and the two-point plane correlator have been used. The integral of $Z_{2}$ is convergent, and we expect that every order of the perturbation expansion of $Z(s)$ is convergent, since the divergent bulk contributions from the numerator $\langle\psi_{0}(0)|\mathcal{T}e^{-\int_{0}^{\beta}\mathrm{d}\tau H_{1}(\tau)}|\mathcal{C}\rangle$ and the denominator $\langle\psi_{0}(0)|\mathcal{T}e^{-\int_{0}^{\beta}\mathrm{d}\tau H_{1}(\tau)}|\psi_{0}(0)\rangle$ cancel at every order. In fact, according to the fusion rule $\varphi_{a}\varphi_{a}\sim 1+\cdots$, we should have $G_{a}(\eta\to 0)\to 1$ and thus $G_{a}(\eta\to 0)-1\to 0$ as $|w_{1}-w_{2}|\to 0$, which cancels the divergent bulk contribution in the integral.
@@
$=2\pi\sum_{n=0}^{\infty}\frac{1}{2n+2h_{a}}\left[\frac{\Gamma(n+2h_{a})}{n!\Gamma(2h_{a})}\right]^{2}\cdot\lim_{\rho\to 0}\left[-\ln\rho-\frac{1-\rho^{2n+2h_{a}}}{2n+2h_{a}}\right]\,,$ (S176)
from which we separated out the second-order contribution of “running” central charge, which is proportional to $-\ln\rho=\frac{2\pi\beta}{L}$, and the second-order contribution of $W(s)$:
@@
$\langle\psi_{0}(s)|\mathcal{C}\rangle$ $=\mathcal{A}_{1}+\left[\frac{\mathcal{A}_{a}}{2(2\pi)^{1-2h_{a}}}\cdot\frac{{}_{2}F_{1}(h_{a},2h_{a},h_{a}+1;-1)}{h_{a}}\right]\cdot s$
@@
$+\frac{\mathcal{A}_{1}}{2(2\pi)^{2-4h_{a}}}\left[\frac{1}{2\pi}\int_{0}^{1}\mathrm{d}\rho_{1}\int_{0}^{1}\mathrm{d}\rho_{2}\int_{0}^{2\pi}\mathrm{d}\theta\frac{(\rho_{1}\rho_{2})^{2h_{a}-1}}{|\rho_{1}-\rho_{2}e^{i\theta}|^{4h_{a}}}\left(G_{a}(\eta)-1\right)-\sum_{n=0}^{\infty}\frac{1}{(2n+2h_{a})^{2}}\left(\frac{\Gamma(n+2h_{a})}{n!\Gamma(2h_{a})}\right)^{2}\right]\cdot s^{2}$
$+\mathcal{O}(s^{3})\,.$ (S178)
@@
The absolute value of the Ishibashi coefficients for the crosscap state $|\mathcal{C}\rangle$ can be determined from the loop channel-tree channel correspondence between the cylinder partition function with two crosscap boundaries and the Klein bottle partition [22]:
$|A_{a}|^{2}=\sum_{b}M_{b,b}S_{b,a}\,,$ (S179)
where $S_{a,b}$ is the modular $\mathcal{S}$ matrix and $M_{a,a}$ is the degeneracy of the diagonal character $\chi_{a}(q)\bar{\chi}_{a}(\bar{q})$ appearing in the modular-invariant partition function on the torus. For many rational CFTs, the data of the matrices $S$ and $M$ are already known, since the modular-invariant partition function can be systematically obtained via the A-D-E classification [62, 63, 12]. Therefore, we can obtain a compact expression for the non-vanishing first-order perturbation correction of the crosscap overlap if $\mathcal{A}_{a}\neq 0$, as shown in Eq. (S178). Specifically, for the $\mathbb{Z}_{3}$ parafermion CFT with thermal perturbation (perturbation operator has conformal weight $h_{\varepsilon}=\bar{h}_{\varepsilon}=2/5$), we have $\mathcal{A}_{1}=(3+6/\sqrt{5})^{1/4}$ and $|\mathcal{A}_{\varepsilon}|=(3-6/\sqrt{5})^{1/4}$, which gives the following first-order perturbation correction to the crosscap overlap:
$\langle\psi_{0}(s)|\mathcal{C}_{\pm}^{\mathbb{Z}_{3}}\rangle=\left(3+\frac{6}{\sqrt{5}}\right)^{1/4}\pm\frac{5}{8(2\pi)^{1/5}}\left(3-\frac{6}{\sqrt{5}}\right)^{1/4}\cdot\frac{\Gamma(2/5)\Gamma(7/5)}{\Gamma(4/5)}\cdot s+\mathcal{O}(s^{2})=1.54401\pm 0.54881\cdot s+\mathcal{O}(s^{2})\,,$ (S180)
Leading order expansion coefficients of the crosscap overlap are expected to be useful in identifying critical theories in lattice models simulations. As an example, we consider the three-state quantum clock chain
with
$\sigma=\begin{pmatrix}0&1&0\\ 0&0&1\\ 1&0&0\end{pmatrix},\quad\tau=\begin{pmatrix}1&0&0\\ 0&e^{2\pi\mathrm{i}/3}&0\\ 0&0&e^{4\pi\mathrm{i}/3}\end{pmatrix},$ (S182)
being $\mathbb{Z}_{3}$ spin matrices, where periodic boundary condition is imposed. At the critical point $h_{3}=1$, the model (S181) is described by the $\mathbb{Z}_{3}$ parafermion CFT, with velocity $v=\frac{3\sqrt{3}}{2}$ determined from the exact solution [64]. Away from (but near) $h_{3}=1$, the model is ordered (with spontaneous $\mathbb{Z}_{3}$ symmetry broken) for $h_{3}<1$ and disordered for $h_{3}>1$, whose low-energy effective theory is just the $\mathbb{Z}_{3}$ parafermion CFT with thermal perturbation. The dimensionless coupling $s$ for the lattice model (S181) is given by $s=\frac{\sqrt{\mathcal{N}_{\varepsilon}}}{v}(h_{3}-1)(vN)^{6/5}$ [40], where $\mathcal{N}_{\varepsilon}=\lim_{r\rightarrow\infty}\lim_{N\rightarrow\infty}r^{8/5}\langle(\tau^{{\dagger}}_{j}+\tau_{j})(\tau^{{\dagger}}_{j+r}+\tau_{j+r})\rangle_{\mathrm{c}}=0.315$ ($\langle\cdots\rangle_{\mathrm{c}}$: connected correlator evaluated at the critical point $h_{3}=1$) is the normalization of the perturbation operator on the lattice. The crosscap overlap between the ground state of Eq. (S181) and the lattice crosscap state $|\mathcal{C}_{\mathrm{latt}}\rangle=\prod_{j=1}^{N/2}\left(\sum_{\alpha=1}^{3}|\alpha\rangle_{j}|\alpha\rangle_{j+N/2}\right)$ has been calculated using DMRG. The numerical results shown in Fig. S1 are in good agreement with the conformal perturbation theory result in Eq. (S180).
Figure S1: Crosscap overlap of the three-state clock model compared with the leading-order (1st order) conformal perturbation prediction. Numerical results with three different chain lengths ($N=20,30,40$) are shown with different colored symbols. For all three data sets, fitting their slopes gives (almost) the same numerical value $\mathcal{C}^{\mathrm{num}}_{1}\approx-0.55$.
F.3 C. Apply to the perturbed Ising CFT
@@
Since the crosscap correlators have been obtained via the bosonziation technique, it is straightforward to apply our result in Eq. (S178) to the Ising field theory, where $|\mathcal{C}_{\pm}\rangle=\mathcal{A}_{1}|1\rangle\rangle_{\mathcal{C}}\pm\mathcal{A}_{\varepsilon}|\varepsilon\rangle\rangle_{\mathcal{C}}$ with $\mathcal{A}_{1}=\sqrt{\frac{2+\sqrt{2}}{2}}$ and $\mathcal{A}_{\varepsilon}=\sqrt{\frac{2-\sqrt{2}}{2}}$.
@@
$\langle\psi_{0}(s_{1})|\mathcal{C}_{\pm}\rangle=\sqrt{\frac{2+\sqrt{2}}{2}}\pm\sqrt{\frac{2-\sqrt{2}}{2}}\cdot\frac{\pi}{4}\cdot s_{1}-\sqrt{\frac{2+\sqrt{2}}{2}}\cdot\frac{\pi^{2}}{32}\cdot s_{1}^{2}+\mathcal{O}(s_{1}^{3})$ (S183)
with $s_{1}=g_{1}L$, which is consistent with the exact solution in Ref. [40].

### v2

@@ Abstract
$\phantom{=}\;\times\prod_{j=1}^{2n}\left[|w_{j}|^{-2m\epsilon_{j}}\left(1+\frac{1}{w_{j}\bar{w}_{j}}\right)^{-1/4}\right]\prod_{1\leq j<l\leq 2n}\left|\frac{1}{w_{j}-w_{l}}\cdot\left(1+\frac{1}{w_{j}\bar{w}_{l}}\right)\right|^{-\epsilon_{j}\epsilon_{l}/2}\,.$ (S169)
@@
${}_{\mathrm{NS}}\langle 0|\sigma(w_{1},\bar{w}_{1})\sigma(w_{2},\bar{w}_{2})|\mathbbm{1}\rangle\rangle_{\mathcal{C}}$ $=\frac{1}{|w_{1}-w_{2}|^{1/4}}\sqrt{\frac{1}{2}\left[(1-\eta)^{-1/4}+(1-\eta)^{1/4}\right]}$
@@
${}_{\mathrm{NS}}\langle 0|\sigma(w_{1},\bar{w}_{1})\sigma(w_{2},\bar{w}_{2})|\mathcal{C}_{\pm}\rangle$ $=\sqrt{\frac{2+\sqrt{2}}{2}}\frac{1}{|w_{1}-w_{2}|^{1/4}}\,(1-\eta)^{-1/8}\left[\frac{\sqrt{2}}{2}\sqrt{1+\sqrt{1-\eta}}\pm\frac{2-\sqrt{2}}{2}\sqrt{1-\sqrt{1-\eta}}\right]\,$, (S173)
where $\eta$ is the crosscap cross ratio [Eq. (S158)]. This corresponds to Eq. (14) in the main text.
@@
The conformal perturbation theory for computing crosscap overlap is given in Eqs. (16) and (17) in the main text. Here, we provide a simple derivation and then formulate the perturbation expansion for a general crosscap state $|\mathcal{C}\rangle$ up to second order. The application to the perturbed Ising CFT is specifically discussed.
@@
where we expect that the perturbed ground-state energy scales as $E_{0}=-\frac{\pi}{6L}c(s)$ and the (finite-size) energy gap scales as $\Delta E(s,L)\sim\frac{1}{L}$, with $s$ fixed. Here, $c(s)$, with $c(0)=c$, is the “running” central charge, denoting the deformed central charge under the perturbation. As $\beta\to\infty$, the contribution of the excited states decays exponentially, leaving only the ground-state contribution. Therefore, in the interaction picture, we can extract $W(s)$ as follows:
[…]
@@
is the change of “running” central charge.
The formal perturbation series can be calculated by expanding the time-ordered exponential $\mathcal{T}e^{-\int_{0}^{\beta}\mathrm{d}\tau\,H_{1}(\tau)}$ in powers of the coupling $g$. Specifically, for $W(s)$, we expand $\langle\psi_{0}(0)|\mathcal{T}e^{-\int_{0}^{\beta}\mathrm{d}\tau\,H_{1}(\tau)}|\psi_{0}(0)\rangle$ as
@@
where each order of the perturbation series can be split into a divergent term $\frac{2\beta}{6L}C_{2n}$, scaled as $\frac{\beta}{L}$ when $\beta\to\infty$, and a subleading finite term $W_{2n}$. The series expansion of $W(s)$ is obtained as $W(s)=\sum_{n=1}^{\infty}W_{2n}s^{2n}$ and the “running” central charge is obtained as $c(s)=c+\sum_{n=1}^{\infty}C_{2n}s^{2n}$.
@@
A crosscap state $|\mathcal{C}\rangle$ can be characterized by the one-point conformal crosscap correlator of primary fields $\varphi_{a}$,
@@
where $\mathcal{A}_{a}$ is the crosscap coefficient. The overall phase of the crosscap state can always be fixed by requiring the coefficient associated with the identity operator to be positive, $\mathcal{A}_{0}>0$. Furthermore, the two-point conformal crosscap correlator can be determined up to an unknown prefactor $G_{a}(\eta)$ as a function of the crosscap cross ratio [22, 33, 34]:
$\langle\psi_{0}(0)|\varphi_{a}(w_{1},\bar{w}_{1})\varphi_{a}(w_{2},\bar{w}_{2})|\mathcal{C}\rangle=\frac{\mathcal{A}_{0}\cdot G_{a}(\eta)}{|w_{1}-w_{2}|^{4h_{a}}}\,$, (S182)
@@
with $Z_{0}=\mathcal{A}_{0}$ by definition. Then, the perturbative expansion of the crosscap overlap up to second order is given by
$\langle\psi_{0}(s)|\mathcal{C}\rangle=\mathcal{A}_{0}+Z_{1}\cdot s+\left(Z_{2}+\frac{1}{2}\mathcal{A}_{0}\,W_{2}\right)\cdot s^{2}+\mathcal{O}(s^{3})\,.$ (S185)
@@
$Z_{1}\cdot s=-\int_{0}^{\infty}\mathrm{d}\tau\langle\psi_{0}(0)|H_{1}(\tau)|\mathcal{C}\rangle=g\int_{0}^{\infty}\mathrm{d}\tau\int_{0}^{L}\mathrm{d}x\,\langle\psi_{0}(0)|\varphi_{a}(z,\bar{z})|\mathcal{C}\rangle\,$, (S186)
@@
$Z_{2}\cdot s^{2}$ $=\frac{1}{2}\int_{0}^{\infty}\mathrm{d}\tau_{1}\int_{0}^{\infty}\mathrm{d}\tau_{2}\left[\langle\psi_{0}(0)|H_{1}(\tau_{1})H_{1}(\tau_{2})|\mathcal{C}\rangle-\mathcal{A}_{0}\langle\psi_{0}(0)|H_{1}(\tau_{1})H_{1}(\tau_{2})|\psi_{0}(0)\rangle\right]$
$=\frac{g^{2}}{2}\int_{0}^{\infty}\mathrm{d}\tau_{1}\int_{0}^{L}\mathrm{d}x_{1}\int_{0}^{\infty}\mathrm{d}\tau_{2}\int_{0}^{L}\mathrm{d}x_{2}\left[\langle\psi_{0}(0)|\varphi_{a}(z_{1},\bar{z}_{1})\varphi_{a}(z_{2},\bar{z}_{2})|\mathcal{C}\rangle-\mathcal{A}_{0}\langle\psi_{0}(0)|\varphi_{a}(z_{1},\bar{z}_{1})\varphi_{a}(z_{2},\bar{z}_{2})|\psi_{0}(0)\rangle\right]\,$, (S187)
@@
$Z_{2}\cdot s^{2}$ $=\frac{g^{2}}{2}\left(\frac{2\pi}{L}\right)^{4h_{a}}\mathcal{A}_{0}\int_{0}^{\infty}\mathrm{d}\tau_{1}\int_{0}^{L}\mathrm{d}x_{1}\int_{0}^{\infty}\mathrm{d}\tau_{2}\int_{0}^{L}\mathrm{d}x_{2}\frac{|w_{1}w_{2}|^{2h_{a}}}{|w_{1}-w_{2}|^{4h_{a}}}\left(G_{a}(\eta)-1\right)$
$=\left[\frac{\mathcal{A}_{0}}{2(2\pi)^{3-4h_{a}}}\int_{0}^{1}\mathrm{d}\rho_{1}\int_{0}^{1}\mathrm{d}\rho_{2}\int_{0}^{2\pi}\mathrm{d}\theta\frac{(\rho_{1}\rho_{2})^{2h_{a}-1}}{|\rho_{1}-\rho_{2}e^{i\theta}|^{4h_{a}}}\left(G_{a}(\eta)-1\right)\right]\cdot s^{2}\,$, (S191)
where the new integral variables read $(\rho_{1},\rho_{2},\theta)=(e^{-\frac{2\pi}{L}\tau_{1}},e^{-\frac{2\pi}{L}\tau_{2}},\frac{2\pi(x_{1}-x_{2})}{L})$. In the above derivation, the one-point and two-point crosscap correlators [Eqs. (S181) and (S182)] and the two-point plane correlator have been used. The integral for $Z_{2}$ is convergent, and we expect that every order of the perturbation expansion of $Z(s)$ is convergent, since the divergent bulk contributions from the numerator $\langle\psi_{0}(0)|\mathcal{T}e^{-\int_{0}^{\beta}\mathrm{d}\tau H_{1}(\tau)}|\mathcal{C}\rangle$ and the denominator $\langle\psi_{0}(0)|\mathcal{T}e^{-\int_{0}^{\beta}\mathrm{d}\tau H_{1}(\tau)}|\psi_{0}(0)\rangle$ cancel at each order. In fact, according to the fusion rule $\varphi_{a}\varphi_{a}\sim\mathbbm{1}+\cdots$, we should have $G_{a}(\eta\to 0)\to 1$ and thus $G_{a}(\eta\to 0)-1\to 0$ as $|w_{1}-w_{2}|\to 0$, which cancels the divergent bulk contribution in the integral.
@@
$=2\pi\sum_{n=0}^{\infty}\frac{1}{2n+2h_{a}}\left[\frac{\Gamma(n+2h_{a})}{n!\Gamma(2h_{a})}\right]^{2}\,\lim_{\rho\to 0}\left[-\ln\rho-\frac{1-\rho^{2n+2h_{a}}}{2n+2h_{a}}\right]\,$, (S195)
from which we separated out the second-order contribution of the “running” central charge, which is proportional to $\ln\rho=\frac{2\pi\beta}{L}$, and the second-order contribution of $W(s)$:
@@
$\langle\psi_{0}(s)|\mathcal{C}\rangle$ $=\mathcal{A}_{0}+\mathcal{A}_{a}\cdot\frac{{}_{2}F_{1}(h_{a},2h_{a},h_{a}+1;-1)}{2(2\pi)^{1-2h_{a}}h_{a}}\cdot s$
@@
$+\frac{\mathcal{A}_{0}}{2(2\pi)^{2-4h_{a}}}\left[\frac{1}{2\pi}\int_{0}^{1}\mathrm{d}\rho_{1}\int_{0}^{1}\mathrm{d}\rho_{2}\int_{0}^{2\pi}\mathrm{d}\theta\frac{(\rho_{1}\rho_{2})^{2h_{a}-1}}{|\rho_{1}-\rho_{2}e^{i\theta}|^{4h_{a}}}\left(G_{a}(\eta)-1\right)-\sum_{n=0}^{\infty}\frac{1}{(2n+2h_{a})^{2}}\left(\frac{\Gamma(n+2h_{a})}{n!\Gamma(2h_{a})}\right)^{2}\right]\cdot s^{2}$
$+\mathcal{O}(s^{3})\,$, (S197)
where $\mathcal{A}_{a}$ is the crosscap coefficient associated with the primary field $\varphi_{a}$, as defined in Eq. (S181).
In rational CFTs, a crosscap state $|\mathcal{C}\rangle$ can be expressed as a linear combination of crosscap Ishibashi states:
@@
where the expansion coefficients are precisely the crosscap coefficients. The absolute values of these coefficients can be determined from the loop channel-tree channel correspondence between the Klein bottle partition function and the cylinder partition function with crosscap boundaries. In particular, for rational CFTs with charge conjugation modular invariants, one can always construct the PSS crosscap state $|\mathcal{C}_{0}\rangle$ associated with the identity field, whose crosscap coefficients are
$\mathcal{A}_{a}=\frac{P_{0a}}{\sqrt{S_{0a}}}\,$, (S199)
as shown in Eq. (S68).
The conformal perturbation theory provides a perturbative approach to investigate the conjectured monotonicity of the Klein bottle entropy (norm-square of the crosscap overlap) under the renormalization flow [43]. Specifically, the first-order correction of the crosscap overlap in Eq. (S197) takes a remarkably simple form. From the series expansion of the hypergeometric function,
$_{2}F_{1}(h_{a},2h_{a},h_{a}+1;-1)=\sum_{n=0}^{\infty}\frac{(2h_{a})_{2n}}{(2n)!}\frac{h_{a}}{2n+h_{a}}\left(1-\frac{2n+2h_{a}}{2n+1+h_{a}}\frac{2n+h_{a}}{2n+1}\right)\,$, (S200)
where $(q)_{n}=q(q+1)\cdots(q+n-1)$ is the Pochhammer symbol, one sees that $_{2}F_{1}(h_{a},2h_{a},h_{a}+1;-1)$ is strictly positive for $h_{a}<1$. Consequently, whenever the first-order correction is non-vanishing, its sign is determined solely by the crosscap coefficient $\mathcal{A}_{a}$.
F.3 C. Perturbed $A$-series unitary minimal models
We apply the conformal perturbation theory to the $A$-series unitary minimal models, with central charge $c=1-\frac{6}{m(m+1)}$, $m\geq 3$. The primary fields are labeled by integer pairs $(r,s)$ within the conformal grid $1\leq r\leq m-1$, $1\leq s\leq m$, subject to the condition that $r+s$ is even to avoid double counting. Since all primary fields are self-conjugate, these theories admit charge conjugation modular invariant torus partition functions.
The PSS crosscap state associated with the identity field $(1,1)$ can be expressed in terms of the Ishibashi states as
$|\mathcal{C}_{(1,1)}\rangle=\sum_{1\leq r\leq m-1\,,1\leq s\leq m}^{r+s=\mathrm{even}}\mathcal{A}_{(r,s)}|(r,s)\rangle\rangle_{\mathcal{C}}\,$, (S201)
where the crosscap coefficients $\mathcal{A}_{(r,s)}$ can be computed via Eq. (S199).
The modular $\mathcal{S}$-matrix is
$S_{(r,s),(r^{\prime},s^{\prime})}$ $=\delta^{(2)}_{r+s,0}\delta^{(2)}_{r^{\prime}+s^{\prime},0}\sqrt{\frac{8}{m(m+1)}}\sin\left(\frac{\pi rr^{\prime}}{m}\right)\sin\left(\frac{\pi ss^{\prime}}{m+1}\right)\,$, (S202)
and the modular $\mathcal{T}$-matrix is determined via the conformal weights of the primary fields
$h_{(r,s)}=\frac{[r(m+1)-sm]^{2}-1}{4m(m+1)}\,.$ (S203)
Using Eq. (S69), one can calculate the associated $P$-matrix as [57, 23],
$\quad P_{(r,s),(r^{\prime},s^{\prime})}$
$=\frac{4}{\sqrt{m(m+1)}}\left[\delta^{(2)}_{r,0}\delta^{(2)}_{s,0}\delta^{(2)}_{r^{\prime},0}\delta^{(2)}_{s^{\prime},0}\sin\left(\frac{\pi rr^{\prime}}{2m}\right)\sin\left(\frac{\pi ss^{\prime}}{2(m+1)}\right)+\delta^{(2)}_{r,1}\delta^{(2)}_{s,1}\delta^{(2)}_{r^{\prime},1}\delta^{(2)}_{s^{\prime},1}\cos\left(\frac{\pi rr^{\prime}}{2m}\right)\cos\left(\frac{\pi ss^{\prime}}{2(m+1)}\right)\right]\,$, (S204)
where $\delta^{(2)}_{n,m}\equiv\frac{1+(-1)^{n+m}}{2}$ is the Kronecker-$\delta$ function with periodic $2$.
Therefore, the crosscap coefficients of the standard PSS state $|\mathcal{C}_{(1,1)}\rangle$ read
$\mathcal{A}_{(r,s)}=\frac{P_{(1,1),(r,s)}}{\sqrt{S_{(1,1),(r,s)}}}=\delta^{(2)}_{r,1}\delta^{(2)}_{s,1}\sqrt{\frac{\sqrt{2}\cot\left(\frac{\pi r}{2m}\right)\cot\left(\frac{\pi s}{2(m+1)}\right)}{\sqrt{m(m+1)}}}\,$, (S205)
which are non-negative for all $(r,s)$. This provides a perturbative proof of the monotonicity conjecture [43] for the crosscap overlap (and hence Klein bottle entropy) under the renormalization group flow in $A$-series unitary minimal models whenever the crosscap coefficient of the perturbing field is non-vanishing.
F.4 D. Ising CFT with thermal and magnetic perturbations
@@
For the Ising field theory, we have already obtained the crosscap correlators via the bosonziation technique. It is straightforward to apply these results in Eq. (S197), where $|\mathcal{C}_{\pm}\rangle=\mathcal{A}_{\mathbbm{1}}|\mathbbm{1}\rangle\rangle_{\mathcal{C}}\pm\mathcal{A}_{\varepsilon}|\varepsilon\rangle\rangle_{\mathcal{C}}$ with $\mathcal{A}_{\mathbbm{1}}=\sqrt{\frac{2+\sqrt{2}}{2}}$ and $\mathcal{A}_{\varepsilon}=\sqrt{\frac{2-\sqrt{2}}{2}}$.
@@
$\langle\psi_{0}(s_{1})|\mathcal{C}_{\pm}\rangle=\sqrt{\frac{2+\sqrt{2}}{2}}\pm\sqrt{\frac{2-\sqrt{2}}{2}}\,\frac{\pi}{4}\cdot s_{1}-\sqrt{\frac{2+\sqrt{2}}{2}}\,\frac{\pi^{2}}{32}\cdot s_{1}^{2}+\mathcal{O}(s_{1}^{3})$ (S206)
with $s_{1}=g_{1}L$. This result is consistent with the exact solution in Eq. (S106), as first reported in Ref. [43].